Daily Report for IODP Expedition 311, October 12, 2005


LOCATION: Site U1328 (CAS-06A)


SCIENCE UPDATE: Due to strong winds and severe ship heave conditions (exceeding 8 m), we had to suspend operations at 03:15 hr to wait for the weather conditions to improve, eventually pulling out of Hole U1238B at 07:45 hr. Up to that point we had cored the top 56 mbsf in Hole U1328B and had taken two PCS pressure cores (U1328B-4P and U1328B-7P). The first PCS core from only 15.0 mbsf, which was targeted to sample several LWD/MWD imaged high electrical resistivity sections identified in Hole U1328A, yielded 22.4 liters of gas and most likely had contained gas hydrate. The second PCS core (U1328B-7P) was still being degassed at the end of the day. After waiting on the weather for 16 hours, the sea conditions had improved to the point to allow drilling and coring operations to continue. We offset the ship by 15 m from location of Hole U1328B and spudded Hole U1328C at 20:30 hr, drilling down to the maximum depth of Hole U1328B (56.5 mbsf) to resume APC coring.